# OmniQuery: Contextually Augmenting Captured Multimodal Memory to Enable Personal Question Answering

### v1

@@ 6.3. Answer Generation
The model analyzes and reasons which captured memories serve as references for the generated answer. These reference memories are also included in the final answer presented to the user. To enhance the reasoning process, OmniQuery leverages chain-of-thought prompting (Wei et al., 2023), ensuring the generation is more accurate and contextually rich. For specific prompts, please refer to Appendix A.3
Figure 7. Four exemplar results of using OmniQuery to answer hybrid personal memory-related questions.
7. RESULTS
To demonstrate OmniQuery’s capabilities in answering users’ personal memory-related questions, we showcase four challenging examples of answering hybrid queries as shown in Figure 7.
7.1. Experiment Setting
OmniQuery was tested on the personal data of one of the researchers from the team, consisting of 2,590 images and videos downloaded from the researcher’s smartphone. The data spans from mid-March to mid-August 2024, covering various activities, trips, and events, including a trip to CHI 2024, a summer trip to Hawaii, as well as captured memories of restaurants, fitness logs, and more.
7.2. Example Walkthrough
In example one, the user asked about the number of cardio sessions completed over the previous two months. OmniQuery successfully retrieved all relevant memories (photos of the stair master machine taken after each session in April and May) and accurately generated the answer (18 sessions), providing every instance as supporting evidence.
In example two, the user sought the name of a Poke restaurant they visited during CHI 2024 in Honolulu. They remembered taking a photo of the meal but forgot the restaurant’s name. OmniQuery successfully retrieved the relevant memories based on the query and provided the correct answer.
In example three, the user asked about the hotel they stayed at during CHI. Although no explicit memories indicate the stay at the Prince Hotel, OmniQuery inferred the hotel from metadata associated with the photos taken at the venue, such as a nighttime photo of a broken takeout bag and a morning view of the marina.
In example four, the user asked about the length of stay during their “second” visit to Hawaii. OmniQuery was able to identify the second trip (the first was the CHI trip) and accurately count the number of stay during the visit.
@@ 8. USER EVALUATION
We conducted a self-guided user evaluation to let participants install and use OmniQuery on their local machine with their local album data. This is to protect participants’ personal data including sensitive and personal identifiable information. In this section, we discuss the detailed process and the evaluation result.
@@ 8.1. Participants
We recruited 10 participants, including seven from our diary study and three additional participants via word-of-mouth. We recruited participants who have basic programming skills so that they were able to install OmniQuery from source on their local machine. They also consented to the whole process, including that their filtered personal data will be processed via an API service. All the 10 participants (4 male, 6 female, age range = 22 to 29, $age_{mean}$ = 25.3, $age_{SD}$ = 2.63) were fluent or native English speaker. Participants rated their frequency of logging their daily lives as ‘Only record essential information’ (1), ‘Regularly log important events and memorable experiences’ (5) and ‘Actively log my daily life’ (4). Each participant is compensated with \$50 for completing this study.
@@ 8.2. Apparatus
Two different systems were implemented in the user evaluation: the OmniQuery pipeline and a baseline system for comparison. The baseline system leverages a conventional RAG architecture, retrieving related memories based solely on vector similarities between the query and the description of the memory. The same base language model and prompt structure as OmniQuery are then used to generate the answer. For detailed implementation of the baseline, please refer to Appendix B.
As shown in Figure 8, in the studies, participants were presented with a single text input box similar to search engine input boxes. After they typed in the question, they would see two answers generated by the two systems in a randomized order. Two rating questions on the answer’s accuracy and completeness were then shown under each answer, from a scale of 1-5.
@@
8.3. Procedure
The user evaluation involved three stages: (1) system setup, (2) data preparation, and (3) the main testing session using user queries.
System setup
Participants were given the source code for OmniQuery to install the back-end and a web application on their local machine. They had the option of an online walkthrough session with the researchers or following the setup instructions on a self-guided manner.
Data preparation
Participants were asked to transfer a set of captured memories (photos and videos) from their smartphone’s album to their laptops since OmniQuery was installed on their local laptops. They were also asked to manually review and filter out any content deemed sensitive or preferred to be excluded from the study. After this filtering process, a total of 1,000 photos and videos were selected for data preprocessing.
Depending on how frequently participants logged their daily lives, the 1,000 selected photos and videos spanned anywhere from one to four months of past memories. These captured memories were then fed into OmniQuery for the query-agnostic augmentation process. The safety of the process data is ensured following the API’s privacy protocol [^6] https://openai.com/policies/privacy-policy/
@@ 8.3. Procedure
The main session lasts 45 minutes, in which participants tested OmniQuery using two types of questions: (1) questions logged during the diary study and (2) questions they generated during the session. Participants checked on the diary study questions manually to determine if they could be answered using the filtered set of data on captured memories. Additionally, participants were encouraged to brainstorm and use new questions that were potentially answerable using the filtered data to comprehensively test OmniQuery.
In the question-answering procedure, OmniQuery-generated answers are accompanied with answers generated from the baseline system implemented using RAG. Each system generated answers anonymously, and participants compared and rated the results for both systems. The answers and user ratings were recorded for quantitative analysis. Throughout the process, participants were asked to think aloud (Nielsen, 1994), and a brief interview was conducted at the end of the session to gather feedback and suggestions. These results were recorded for qualitative analysis.
@@ 8.4. Comparison Metrics
After two answers were presented for a question, participants were asked to rate the two answers. We used the Chatbot Arena evaluation method (Chiang et al., 2024), where participants compared answers from two systems and selected the better one or marked it as a tie. More specifically, for each question, participants rated the user perceived accuracy (UPA) and user perceived completeness (UPC) of the answers from both systems.
The UPA score was rated on a scale from 1 to 5: 1: Completely wrong or invalid result; 2: Incorrect, but provides at least some insight that helps answer or further refine the question; 3: Partially correct, or contains a subset of correct answers (e.g., only listing one meal when asked about all meals eaten last week); 4: Mostly correct, but missing some minor details (e.g., missing one subway trip when asked how many times I rode the subway); and 5: Completely correct. The UPC score, on the other hand, focused on the completeness and credibility of the given answers. In most cases, the questions asked by participants were challenging and needed to be explained and powered by evidence from the captured memory data. For example, when asked questions which have numeric answers like “how many meals did I have during my last New York trip,” the systems could be right on the numeric answer but wrong at which meals were counted. In those cases, participants examined the answers by looking back at the filtered data, and gave ratings on how they feel about the completeness and credibility of each answer.
We also directly compare the ratings of OmniQuery and the baseline to determine if one can outperform the other. In the comparison, if both systems have a UPA of 2 or lower (incorrect), the result is labeled ”both are bad.” If at least one system provides a better-than-”partially correct” answer ($\geq$3), the system with the higher UPA is considered the winner. In cases where both systems have the same UPA, the one with the higher UPC is the winner. Otherwise, it is a tie.
@@ 8.5. Quantitative Result
Participants tested 137 queries in total during the main session. Among them, 28 were previously logged during the diary study. We manually labeled each tested query using the categorization and definition mentioned in section 3.3. As a result, 24 were categorized as direct content query while 17 were contextual filters and 96 were hybrid queries. We analyzed the performance metrics of both systems (OmniQuery and baseline) using the scores rated by the participants. Table 2 and 3 summarize our results. In addition to presenting the average UPA and UPC scores, we calculated binary accuracy to evaluate whether the systems provided mostly correct answers. An answer was considered accurate if its UPA score was equal to or greater than 4 (mostly correct) (Table 2). We also present the “comparison result” in Table 3, which compares the two systems head-to-head on answering personal questions.
The result shows that overall, OmniQuery outperforms the baseline system in both the accuracy and completeness. Specifically, OmniQuery achieves an accuracy of 71.5%, outperforming the baseline by 28.4%, winning the comparison 52.6% of the time, and tying 21.9% of the time, while in 14.6% of the time, both results are bad. We also present the results for different categories of queries. The results indicate that simpler techniques like the baseline handle direct content queries reasonably well (62.5 % accuracy, and winning or tying 41.6% of the time). While the baseline struggles with more complex queries such as contextual filters or hybrid queries (38.9% accuracy, winning or tying 31.0% of the time), OmniQuery demonstrates it capabilities in effectively handling such queries (69.0% accuracy, winning or tying 72.6% of the time).
@@ 8.6. Qualitative Feedback and Findings
Apart from quantitative analysis, we also present qualitative feedback and findings from the think-aloud protocol and the exit interview on the usage of OmniQuery and suggestions of such intelligent question-answering system.
8.6.1. Comparing with Existing Tools
All participants have tried searching objects in their smartphone albums in their daily lives. They reported that they used the search features mostly to find a specific piece of information, like driver license, SSN number on the card, etc., matching the first type of questions (Direct Content Query) listed in section 3.3. They would also search for a specific event, like a trip to a specific location, matching the second type of questions (Contextual Filter). However, these searches are usually limited to retrieving a clear and specific object that users were looking for. They could not handle more complex questions like were collected in our diary studies. Plus, some of our participants also anticipated for this to happen because they “know what can be searched and what cannot be searched” from these existing album search tools (P2).
In the studies, a lot more challenging questions were asked. For Direct Content Queries, it would be challenging to answer when the object is ambiguous or when the users can only describe the object and do not know the exact name of it. For Contextual Filter and Hybrid or even more open-ended and subjective questions, existing searching tools are not comparable to OmniQuery and the baseline at all because tools like iOS album search only return specific photos and videos in a whole without contextual understanding or filtering. In other words, these tools are all not “intelligent enough” to our participants.
In this subsection, we further summarize the cases that are hard to be accomplished using existing tools. In comparison to the high-level question types provided in section 3.3, we dived deeper into what these questions in the study were about, and provided detailed examples of these cases.
@@ 8.6.1. Comparing with Existing Tools
  • Look up and Locate: When users know specific references or attributes about an item, such as date, location, or a person in the photo, and want to quickly locate the relevant media. For example, P4 asked, “Can you find the photo of me on a flyer on Instagram?”
@@
8.6.2. Reaction to Answers
Participants reacted differently to different types of answers. They provided their observations as well as perceived feelings on the answers. Note that since participants weren’t aware of which answer was provided by OmniQuery or the baseline RAG system, we present their feedback on the overall answer structure, given that both OmniQuery and the baseline system provided a similar structure of answer and supporting materials.
Dynamics of Detail and Concise Answer. P7 mentioned that they did not always prefer to have a detailed answer composing all related media materials. More reference media would reduce the credibility of the answer to them. P8 also pointed out that if the answer just included all media it could find without a clear and concise connection to the answer, he “might just use iOS album search to get all photos containing [a specific object] and look by myself.”
Failure Cases of Answers. We also present cases when participants reacted negatively to the answers. All participants encountered cases where the answers are inaccurate. Some were incomplete (e.g., P1 believed that they visited mroe than a few churches on the trip to Barcelona, but answers provided only two of them). Some were presumptive (e.g., P7 asked about recent social events, where the answers gave a piece of memory on a museum visit and explained that visiting museums is “likely with other people.” However, P7 visited the museum alone). Some were making mistakes (e.g., P7 asked for the mostly visited attractions but both system mistakenly answered a museum, which was because P7 took a lot of museum pictures and both systems failed to recognize that they were the same visit.) Some even more challenging questions that caused failure of both systems include questions relate to a specific person. For example, P8 asked about her significant other and P5 asked about their “Korean friend” met in a trip. These cases represent the difficulty of understanding the nuances of personal relationships with personal album data.
8.6.3. Iterative Editing of Questions
Another recurring theme in our study is that six participants mentioned the possibility of iterating on the questions based on the answers. It occurred when participants were uncertain of what answers they were gonna get out of open-ended or subjective answers. They wanted to iterate on the questions based on the given answers, which gave them more understanding on the questions they wanted to ask. For example, P3 asked about how many places they visited during summer, and, based on the answer, realized they were more interested in the number of cities rather than the countries. It highlighted the potential need of a chat interface of personal captured memory data, which posit more challenges in this domain considering the query histories.
@@ 9. DISCUSSION
In this section, we draw on implications from our studies to discuss limitations and propose future work based on what we found.
@@ 9.1. From Chat Interface to Multimodal Interactions
As a system designed to answer users’ questions on their personal captured memory, OmniQuery is currently designed in an ask-and-react manner for the purpose of evaluating its efficacy in a lab-study settings. In our studies, participants were excited about what OmniQuery was capable of, and gave feedback on having more multimodal interactions rather than a “ChatGPT-like” interface. We recognize the potential of a more interactive OmniQuery in the following ways:
Multimodal Input and Output: Just like how ChatGPT iterated, OmniQuery could use more multimodal input including but not limited to audio, image, and even videos in the future. Recalling the summarized task cases which can be hard for existing album searching tools to accomplish, a number of these cases can inherently benefit from multimodal inputs. For example, users might want to look up and locate an oddly-shaped cup that they cannot refer to in plain text, or compare all existing dresses in album with a reference color on an image that is hard to describe. Natural language description’s limitation needs to be addressed by introducing different input modalities. Besides input, OmniQuery also has huge potential in helping users relive their memories by visualizing their captured memory data. Users can access their captured memories in an interactive way, inspect and even modify annotations in a transparent way. With that, a “mind palace” style AI assistant becomes more possible.
Error correction: In our studies, we observed the importance of allowing users to review and access the identified composite contexts and semantic knowledge. Participants expressed the need to correct errors within the system when it retrieved incorrect or irrelevant information. For example, P9 asked about a specific KPop store, and while the system successfully generated relevant memories, it mistakenly included an Instagram screenshot of a Korean TV show. In such cases, participants wanted the ability to mark these irrelevant results as incorrect, suggesting that incorporating error correction mechanisms would enhance the system’s overall performance and accuracy in future interactions.
Follow-up queries: Another recurring theme in our study is that participants frequently desired the ability to refine their queries or ask follow-up questions. Six out of ten participants mentioned the need to clarify answers or iteratively narrow down their queries. For instance, P4 and P8 emphasized the importance of follow-up questions to hone in on specific information after receiving initial responses. Together with proposed multimodal interactions above, an iterative interaction model, similar to a chatbot workflow, would significantly improve usability and flexibility of OmniQuery.

### v2

@@ 6.3. Answer Generation
The model analyzes and reasons which captured memories serve as references for the generated answer. These reference memories are also included in the final answer presented to the user. To enhance the reasoning process, OmniQuery leverages chain-of-thought prompting (Wei et al., 2023), ensuring the generation is more accurate and contextually rich (specific prompts in Appendix A.4).
Figure 7. For each captured memory, the baseline RAG system lacks the contextual information extracted through the taxonomy-based augmentation used in OmniQuery.
@@ 7. USER EVALUATION
We conducted a user evaluation to test OmniQuery’s capabilities in handling real-world personal data by comparing it against a baseline system implemented with a typical RAG structure for question answering. Both systems were deployed on the participants’ local machine to protect their personal data. In this section, we discuss the detailed evaluation process, metrics, and results, including quantitative results of the two systems’ performances, representative examples, and qualitative feedback.
@@ 7.1. Participants
We recruited 10 participants, including seven from our diary study and three additional participants via word-of-mouth. They consented to the whole process, including that their filtered personal data will be processed via an API service. All 10 participants four male, six female, age range 22 - 29, $\bar{x}=25.3$, $\mathrm{SD}=2.63$) were fluent or native English speakers. The participants rated their frequency of logging their daily lives as ‘Only record essential information’ (1), ‘Regularly log important events and memorable experiences’ (5) and ‘Actively log my daily life’ (4). Each participant was compensated with \$50 for completing this study.
@@ 7.2. Apparatus
Two different systems were implemented in the user evaluation: the OmniQuery pipeline and a baseline system for comparison. The baseline RAG is designed to differ only in OmniQuery’s core contributions. As shown in Figure 7, while the baseline RAG includes basic contextual information for each captured memory, such as time, location, and a detailed description of the scene, the key differences are: (1) The contextual information remains “raw” and is not structured using the taxonomy derived in OmniQuery, and (2) it lacks composite context and semantic knowledge, as it does not extract contextual information from multiple related memories. Additionally, in the question-answering phase, the baseline does not leverage taxonomy-based retrieval. All other components, such as the base LLM and prompt structure, remain the same. This design ensures the delta between OmniQuery and the baseline highlights and evaluates OmniQuery’s core contributions: taxonomy-based augmentation and retrieval. For the detailed implementation of the baseline, please refer to Appendix B.
As shown in Figure 8, in the studies, participants were presented with a single text input box similar to search engine input boxes. After they typed in the question, they would see two answers generated by the two systems in a randomized order. Two rating questions on the answer’s accuracy and completeness were then shown under each answer with a scale of 1-5.
@@
7.3. Procedure
The user evaluation has three stages: (1) system setup on participants’ local machines, (2) personal data preparation, and (3) the main testing session.
System setup
The participants were given the source code for OmniQuery to install the back-end and a web application on their local machine. They had the option of an online walkthrough session with the experimenters or following the setup instructions on a self-guided manner.
Personal data preparation
@@ 7.3. Procedure
To test OmniQuery on local machines, the participants were asked to transfer a set of captured memories (both photos and videos) from their smartphones’ albums to their laptops. To further protect participants’ privacy, they were instructed to manually review and filter out any content deemed sensitive or preferred to be excluded from the study. This process needs to meet two key requirements: (1) the manual filtering should not become an excessive burden for participants, and (2) the transferred data should be sufficiently large to simulate real-world usage, containing diverse contexts of distant memories. It is impractical to include all participants’ memories, which average 13630 files as collected from the diary study. Thus we perform the following calculation to balance the two requirements:
$T\geq\frac{C}{K}\times\frac{1+R}{1+F\times R}$
Where $T$ is the total number of files, $C$ the context limit, $K$ the token cost per frame, $R$ the relevant frame ratio, and $F$ the average frames per video. The equation balances manual filtering and data diversity. The rationale behind the equation is that processing all memories at each query should exceed the limit of existing powerful models, and thus motivating the need for accurate retrieval to answer queries. Specifically, the model OmniQuery uses has a context window of 128K [^7] https://platform.openai.com/docs/models#gpt-4o, and the minimum token cost for processing each frame when employing low-fidelity understanding is 85888https://platform.openai.com/docs/guides/vision. Ratio represents the video-to-image ratio, which averages 0.12 ($\mathrm{SD}=0.11$) based on our diary study survey. As discussed in Sec. 5.4, the number of frames sampled per video during processing is 10. This calculation results in a total number of files of ~767. To expand the memory coverage within the participants’ manual filtering capacity, we round this up to 1,000 files.
Depending on how frequently participants logged their daily lives, the memory coverage of the 1,000 selected photos and videos spanned from one to four months ($\bar{x}=2.3$, $\mathrm{SD}=1.03$). While this coverage is smaller than what a smartphone album can cover in real-world scenarios (several years), it still represents a moderately distant memory range that is not too recent to demonstrate the capabilities of OmniQuery. These captured memories were then fed into OmniQuery for the taxonomy-based query-agnostic augmentation process. The safety of the process data is ensured following the API’s privacy protocol [^9] https://openai.com/policies/privacy-policy/.
@@
The main session lasts 45 minutes, in which the participants tested OmniQuery using two types of questions: (1) questions logged during the diary study and (2) questions they generated during the session. The participants checked on the diary study questions manually to determine if they could be answered using the filtered set of data on captured memories. Additionally, the participants were encouraged to brainstorm and use new questions that were potentially answerable using the filtered data to comprehensively test OmniQuery. Note that participants have no access to the contextual information augmented by OmniQuery. Instead, they follow a simple mental workflow: attempting to recall, asking the system, and verifying the answer.
In the question-answering procedure, OmniQuery-generated answers are accompanied with answers generated from the baseline system implemented using RAG. Each system generated answers anonymously, and the participants compared and rated the results for both systems. The answers and user ratings were recorded for quantitative analysis. Throughout the process, the participants were asked to think aloud (Nielsen, 1994), and a brief interview was conducted at the end of the session to gather feedback and suggestions. These results were recorded for qualitative analysis.
@@ 7.4. Comparison Metrics
After two answers were presented for a question, the participants were asked to rate the two answers. We used the Chatbot Arena evaluation method (Chiang et al., 2024), where each participant compared answers from the two systems and selected the better one or marked it as a tie. More specifically, for each question, each participant rated the user perceived accuracy (UPA) and user perceived completeness (UPC) of the answers from both systems.
The UPA score was rated on a scale from 1 to 5: 1: Completely wrong or invalid result; 2: Incorrect, but provides at least some insight that helps answer or further refine the question; 3: Partially correct, or contains a subset of correct answers (e.g., only listing one meal when asked about all meals eaten last week); 4: Mostly correct, but missing some minor details (e.g., missing one subway trip when asked how many times I rode the subway); and 5: Completely correct. The UPC score focused on the completeness and credibility of answers, essential for assessing reliability. Many participant queries were complex, requiring explanations supported by captured memory data. For example, when asked “How many meals did I have during my last New York trip?” a system might get the number right but miscount specific meals. Participants reviewed filtered data and rated answers based on their perceived completeness and credibility.
Figure 9. Four representative examples of using OmniQuery to answer hybrid personal memory-related questions.
We also directly compare OmniQuery and the baseline to analyze performance. If both have a UPA of 2 or lower, the result is labeled “both are bad”. If at least one scores $\geq$3, the system with the higher UPA wins. In case of a tie on UPA, the system with the higher UPC wins; otherwise, it remains a tie.
8. Results
We report the results of our user evaluation by presenting quantitative results of comparing OmniQuery with the baseline, representative examples from the study, and insights gathered from the think-aloud protocol and exit interviews.
@@ 8.1. Quantitative Result
The participants tested 137 queries in total during the main session. Among them, 28 were previously logged during the diary study. We manually labeled each tested query using the categorization and definition mentioned in section 3.3. As a result, 24 were categorized as direct content query while 17 were contextual filters and 96 were hybrid queries. We analyzed the performance metrics of both systems (OmniQuery and baseline) using the scores rated by the participants. Table 2 and 3 summarize our results. In addition to presenting the average UPA and UPC scores, we calculated binary accuracy to evaluate whether the systems provided mostly correct answers. An answer was considered accurate if its UPA score was equal to or greater than 4 (mostly correct) (Table 2). We also present the “comparison result” in Table 3, which compares the two systems head-to-head on answering personal questions.
The result shows that, overall, OmniQuery outperforms the baseline system in both the accuracy and completeness. Specifically, OmniQuery achieves an accuracy of 71.5%, outperforming the baseline by 28.4%, winning the comparison 52.6% of the time, and tying 21.9% of the time. For 14.6% of the time, both results are bad. We also present the results for different categories of queries. The results indicate that simpler techniques like the baseline handle direct content queries reasonably well (62.5 % accuracy, and winning or tying 41.6% of the time). While the baseline struggles with more complex queries such as contextual filters or hybrid queries (38.9% accuracy, winning or tying 31.0% of the time), OmniQuery demonstrates it capabilities in effectively handling such queries (69.0% accuracy, winning or tying 72.6% of the time). Specifically, for the queries logged during the diary study, OmniQuery achieved results similar to its overall performance (67.9% accuracy, and winning or typing 74.4% of the time).
8.2. Representative Examples
We selected four representative examples tested by the participants in the evaluation, which are illustrated in Figure 9.
  1. (1)
    P3 wanted to recall the name of the whiskey they tasted during their trip, as they enjoyed it and wanted to check the price. OmniQuery successfully retrieved the target memory (a photo of the four bottles) and generated a specific answer about the bottles they might have ordered.
  2. (2)
    P6 wanted to organize the footage from their visit to a farm and asked about the first animal they petted there. OmniQuery accurately retrieved the memories related to the composite context (“visit to the far”) and used the temporal order to generate the correct answer: the first animal petted was a sheep.
  3. (3)
    P1 wanted to estimate how many times they had practiced golf in the past month to track their progress. OmniQuery successfully retrieved all relevant memories, including videos of golf swings at the driving range and photos taken at Topgolf, and accurately generated the answer.
  4. (4)
    P9 wanted to recall the name of the lake where they went paddling with friends. While no direct memory of paddling is captured, there are several related photos available, including a paddleboard being pumped next to a lake. OmniQuery successfully retrieved these memories and generated the answer using the metadata associated with them.
@@ 8.3. Qualitative Feedback and Findings
All participants had experience using smartphone album search features, primarily for retrieving specific information like driver’s licenses or events such as trips, aligning with direct content queries and contextual filters (Section 3.3). However, they noted that existing tools are limited to finding specific objects and cannot handle more complex queries. Plus, some of our participants also anticipated for this to happen because they “know what can be searched and what cannot be searched” from these existing album search tools (P2).
In the studies, a lot more challenging questions were asked. For direct content queries, it would be challenging to answer when the object is ambiguous or when the users can only describe the object and do not know its exact name. For contextual filter and hybrid or even more open-ended and subjective questions, existing searching tools are not comparable to OmniQuery and the baseline at all because tools like iOS album search only return specific photos and videos without contextual understanding or filtering.
Here we further summarize the cases that are hard to be accomplished by existing tools. In comparison to the high-level question types provided in Section 3.3, we dive deeper into what these questions in the study were about and provide detailed examples.
@@
  • Look up and Locate: When users know specific references or attributes about an item, such as date, location, or a person in the photo, and want to quickly locate the relevant media, such as “Can you find the photo of me on a flyer on Instagram? (P4)”
@@
8.4. Failure Cases
Among the 137 queries tested, 25 failed due to ambiguity, missing contextual cues, information loss, retrieval redundancy, subjectivity, or unavailable memories. Please refer to Appendix C for detailed discussions.
@@ 9. DISCUSSION
In this section, we draw on implications from our studies to both discuss limitations and propose future work.
9.1. Curating Fixed Dataset for Benchmarking
Benchmarking is a widely used approach to evaluate whether new systems and algorithms achieve state-of-the-art performance on specific tasks. It also enables ablation studies to assess the effectiveness of individual components of the proposed design. Currently, OmniQuery is evaluated as an integrated system to compare against another system to demonstrate the effectiveness of its overall design. As a future direction, curating a fixed benchmark dataset would allow for more granular evaluation of OmniQuery’s performance on personal question answering over multimodal captured memories. This would enable deeper insights into how different design choices (both high-level and low-level) affect task performance through objective ratings. Additionally, it could enable testing multiple system parameters (e.g., sliding window sizes, top-K values for retrieval, or prompt designs for inferring contexts) to achieve the optimal performance. Furthermore, it would also allow ablation studies to assess the impact of individual components (e.g., query augmentation). We further discuss significant challenges of curating a fixed benchmark dataset for this personal data task in Appendix D.
@@ 9.2. From Chat Interface to Multimodal Interactions
As a system designed to answer user queries on their personal captured memory, OmniQuery is currently designed in an ask-and-react manner to evaluate its efficacy in a lab-study setting. In our studies, the participants were excited about what OmniQuery was capable of and gave feedback on having more multimodal interactions rather than just a chat interface. We recognize the potential of a more interactive OmniQuery in the following ways:
Multimodal Input and Output. OmniQuery could support multimodal inputs, including audio, images, and videos, to address limitations of text-based search. Many challenging cases for existing album search tools could benefit from this, such as locating an oddly shaped cup or matching dresses by color. Beyond retrieval, OmniQuery could help users relive memories by visualizing captured data, enabling interactive exploration and annotation edits. This brings us closer to a “mind palace” style AI assistant.
Error correction. In our studies, we observed the importance of enabling users to review and refine identified composite contexts and semantic knowledge. Participants expressed the need to correct errors when the system retrieved irrelevant information. For example, P9 asked about a K-pop store, but the system mistakenly included an Instagram screenshot of a Korean TV show. To address this, we propose integrating error correction mechanisms with explanatory insights, confidence levels, and a verification loop, allowing users to mark errors, refine results, and enhance system accuracy over time.
Follow-up queries. A key theme in our study is participants’ need to refine queries or ask follow-up questions, with six out of ten mentioning the desire to clarify responses or narrow their searches iteratively. This was particularly relevant when errors were perceived, as discussed in Error Correction. To address this, we propose augmenting follow-up interactions with explanations and confidence levels to highlight uncertainties. A top-K retrieval strategy could also provide ranked answers for ambiguous queries, enabling iterative refinement. Future work could evaluate these approaches through a longitudinal study.